OpenStudy (anonymous):

An object moves along the x axis according to the equation x = 3.50t2 − 2.00t + 3.00, where x is in meters and t is in second Determine average acc/ average speed between 1.9s and 3.30s find instantaneous speed / acc at t=1.9 Determine the instantaneous speed / acc at t = 3.30 s. (e) At what time is the object at rest?

OpenStudy (anonymous):

hi you can use from this\[V=\Delta x/\Delta t\] for average v & differentiate respect to t from x equation yields v equation if put t in it instantaneous v yields

OpenStudy (anonymous):

for acc differentiate proportion to t from v equation and put t in it for instantaneous a

OpenStudy (anonymous):

for average a use it\[a=\Delta V/\Delta t\]

OpenStudy (anonymous):

when body recieve to rest it's velosity be zero so find time of that from velocity equation \[0=at+V _{0}\]

OpenStudy (ash2326):

We have \[x=3.5t^2-2.00 t+3.00\] let's find velocity, we know that velocity is given as \(\frac{dx}{dt}\) let's differentiate x with respect to t \[\frac{dx}{dt}= 2*3.5t^{2-1}-2.00+0\] so velocity v \[v=7t-2\] we have to find average speed between 1.9 s and 3.3 seconds so average speed =\( \frac{1}{3.3-1.9}\int_{1.9}^{3.3} (7t-2) dt \) \[ \frac{1}{1.4}[ 7t^2/2-2t]_{1.9}^{3.3}\] \[ \frac{1}{1.4}[(7*3.3^2/2-2*3.3)-(7*1.9^2/2-2*1.9)]\] we get \[\frac{1}{1.4} * 22.68\]= 16.2 m/s

OpenStudy (ash2326):

for instantaneous speed at t=1.9 and 3.3 sec substitute t=1.9 and 3.3 in v=7t-2 for acceleration, differentiate v with respect to time will get a=7 m/s^2 so a is 7m/s^2 at t=1.9 sec and 3.3 sec